40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2382     Entry time: Fri Aug 2 10:23:48 2019
Author: ScottA 
Type: DailyProgress 
Category: RFAM 
Subject: Difference in Transfer function of Bias-tee plus EOM with Different input power 

I replaced the EOM driver that was attached to the south EOM with the Bias Tee and the driver which I tuned with the Bias Tee and Dummy EOM. After some minor tuning, I was able to move the transfer functions peak back over 37 MHz. I took the spectrum of the RF monitoring port while a 16.5 dBm signal was present on the RF in port from OCXO after placing an 8 dBm attenuator. We noticed that the max peak at 37 MHz was around -3.49 dBm which corresponds to a signal with an amplitude of 6.4122 V by converting to voltage, multiplying by sqrt(2) and dividing by the value of the simulated TF at 37MHz which is about 0.0328. This is less than the expected 20 V amplitude from a 16.5 dBm signal with the transfer function from the previous post (https://nodus.ligo.caltech.edu:8081/CTN/2376). This motivated us to take transfer functions at different input powers to see if it does change. We can see that it does a fair amount for large input signals. I can offer no explanation at this time.

Attachment 1: Transfer_EOM_driver_W_Bias_and_REAL_EOM_37.pdf  45 kB  Uploaded Fri Aug 2 11:26:04 2019  | Hide | Hide all
Transfer_EOM_driver_W_Bias_and_REAL_EOM_37.pdf Transfer_EOM_driver_W_Bias_and_REAL_EOM_37.pdf
Attachment 2: REAL_EOM_TF_Spectrums_and_data.zip  227 kB  Uploaded Fri Aug 2 11:28:54 2019
Attachment 3: Bias-Tee-and-Driver.jpg  4.094 MB  Uploaded Fri Aug 2 11:30:03 2019  | Hide | Hide all
Bias-Tee-and-Driver.jpg
ELOG V3.1.3-